OpenStudy (anonymous):

Which of the following are solutions to the equation below 7x^2+35x-252=0 A.4 B.-7 C.-4 D.252 E.-9 F.9

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

You have a choice: You could actually solve the given equation, or you could substitute the possible answers one by one to determine whether any of them is a solution, or you could graph the given equation and determine the roots that way.

OpenStudy (anonymous):

ok ,can u help me

OpenStudy (mathmale):

sure, but I will need and want your participation. I've given you several choices; I'm waiting for you to choose one of the actions I've suggested and to get started doing it.

OpenStudy (anonymous):

subsitute the the possible answers

OpenStudy (cwrw238):

to make it easier you can divide the equation by 7 it then becomes x^2 + 5x - 36 = 0 - you can solve this by factoring

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

you 2 numbers whose product is -36 and whose sum is +5

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

OK, @Twilliams250 : You have chosen to do this problem by substituting the given answers. In this case, why don't we start with x=4, substitute that value, and determine whether the given equation is true or false? 7x^2+35x-252=0 requires that we let x=4 in the first two terms. We then get 7(4)^2 + 35(4) -252. Is that = to 0 or not equal to 0?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

equal

OpenStudy (anonymous):

the question says select all corect answers

OpenStudy (mathmale):

You could also use the approach suggested by cwrw238: to make it easier you can divide the equation by 7 and obtain 7[(4)^2+5(4)-36)=0, which is easier to evaluate. If you find that this equation is TRUE when x=4, then you have found one of the possible correct answers. Just move on to the next given possible answer; substitute this value into 7(x^2+5x-36)=0 and determine whether this eq'n is true or false now. And so on. ...

OpenStudy (anonymous):

ok can u help me get all the correct answers

OpenStudy (anonymous):

so 4 works